Eight blue​ socks, four white​ socks, and six gray socks are mixed in a drawer. You pull out two​ socks, one at a​ time, without looking. Complete parts​ (a) through​ (d). What is the probability of the event of getting 2 socks of the same​ color?
Answer by math_tutor2020(3816)


Answer: 49/153

Work Shown

8 blue, 4 white, 6 gray
8+4+6 = 18 socks total

A = probability of getting 2 blue socks
A = (probability 1st sock is blue)*(probability 2nd sock is blue given 1st is blue)
A = (8/18)*(7/17)
A = 56/306
I'm assuming replacement is not applied.

B = probability of getting 2 white socks
B = (4/18)*(3/17)
B = 12/306

C = probability of getting 2 gray socks
C = (6/18)*(5/17)
C = 30/306

D = probability of getting 2 socks of the same color
D = probability that event A or B or C occurs
D = A+B+C because events A,B,C are pairwise mutually exclusive
D = 56/306 + 12/306 + 30/306
D = (56+12+30)/306
D = 98/306
D = (49*2)/(153*2)
D = 49/153

49/153 = 0.32026 = 32.026% approximately